If you wish to save $100,000 at the end of 5 years, how much will you need to invest each month if the interest rate is 12% compounded monthly?


1
Expert's answer
2021-06-09T13:30:49-0400

Future value of annuity"=p\times \frac{(1+r)^{n}-1}{r}"

 where P is the monthly payment amount. r is equal to the interest rate. n is the number of payments

"r=\frac{12}{100}=0.12\\=\frac{0.12}{12}=0.01"


"100000=p\times\frac{(1+0.01)^{60}-1}{0.01}"


"100000=81.67p\\p=\$1224.44"